OpenStudy (anonymous):

Which system has a solution in which the x value is greater than 1? A coordinate plane showing the graphs of the lines y equals 3 x minus 3, y equals negative one-fourth x plus one, y = 3x+2, and y equals negative two-thirds x A. y = 3x + 2 and y = 3x - 3 B. y = -2/3 x and y = -1/4 x + 1 C. y = 3x + 2 and y = -1/4 x + 1 D. y = 3x - 3 and y = -1/4 x + 1

OpenStudy (vishal_kothari):

A) y = 3x+2 = 3x-3 .................... 2=-3 <-- no solutions B) y = -2/3 x = -1/4x +1 .... (1/4-2/3) x = 1 ......... -5/12 x = 1 ................. x = -12/5 <--- Less than 1 C) y = 3x+2 = -1/4x +1 ...... (3+1/4) x = -1 ........ (13/4) x = -1 ................. x = -4/13 <--- Less than 1 D) y = 3x -3 = -1/4 x + 1 ...... (3+1/4) x = 4 ........ (13/4) x = 4 ................. x = 16/13 <-- greater than 1..

OpenStudy (vishal_kothari):

is it correct?

OpenStudy (anonymous):

yes

OpenStudy (vishal_kothari):

thanks...